OpenStudy (anonymous):

How many solutions can be found for the equation 3x − 2 = 3x + 5? Zero One Two Infinitely Many

OpenStudy (anonymous):

@aryandecoolest @AaronAndyson @autumnm @Adi3 @HELPMEPLZ!! @HelpOfTheGods @CallMeKiki @courtneygraley009 @DarkMoonZ @Loser66

OpenStudy (anonymous):

@michayla223 @GreatGuy @minisweet4 @baconbaby1334 @selfconfidence777

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

What is the value of s in the equation 4(2s − 1) = 7s + 12? 12 16 32 34

OpenStudy (anonymous):

website it says no solution

OpenStudy (anonymous):

ppz help me on this one to

OpenStudy (anonymous):

The steps below show the incomplete solution to find the value of x for the equation 7x − 3x − 4 = −1 + 21: Step 1: 7x − 3x − 4 = −1 + 21 Step 2: 7x − 3x − 4 = 20 Step 3: 4x − 4 = 20 Which of these is most likely the next step? 4x = 24 4x = 16 4x = 5 4x = 80

OpenStudy (anonymous):

@faithfloridia1 @sleepyjess @xlovelystonerx @nerd_3.14

OpenStudy (anonymous):

4x=24